OpenStudy (anonymous):

write in factored form: 1: (cos^2x)+(2cosx+1) 2. 1-2sinx+(1-cos^2x) 3. cosx-2sin^2x+1 4. 4tan^2x-[4/(cotx)]+sinx(cscx) Write each expression as an algebraic expression of a single trigonometric function. 1. (1-sin^2x)/(1+sinx) 2. sin^2x/(1+cosx)

OpenStudy (mathmate):

I'll do the first two: 1. cos^2(x) + 2cos(x) + 1 happens to be a perfect square in cos(x), so =(cos(x)+1)^2 2. 1-2sin(x)+(1-cos^2(x)) 1-2sin(x)+sin^2(x) Again, it's a perfect square, so (1-sin(x))^2 The third one is similar to the first two, and the fourth one, convert sin(x) csc(x) = sin(x)/cos(x) = tan(x) and note that 1/cot(x) = tan(x)

OpenStudy (anonymous):

for 2 i have 2-2sinx-cos^2x

OpenStudy (mathmate):

For 2, you do not want to introduce cos(x), use (1-cos^2(x))=sin^2(x), then you can factorize.

OpenStudy (anonymous):

the third one, also, my answer is that it is not factorable

OpenStudy (mathmate):

What did you have?

OpenStudy (mathmate):

cosx-2sin^2x+1 =cosx-2(1-cos^2(x))+1 =2cos^2(x)+cos(x)-1 =(2cos(x)-1)(cos(x)+1)

OpenStudy (anonymous):

i am very confused :( I am just not getting this chapter

OpenStudy (mathmate):

For this one, you need to know that sin^2(x)+cos^2(x)=1, which allows you to convert sin^2(x) to cos^2(x) or vice versa. Once you have them all in (either) sin(x) or cos(x), it's just like a variable called s or c. So #3 can be looked at as: cosx-2sin^2x+1 =cosx-2(1-cos^2(x))+1 =2cos^2(x)+cos(x)-1 =2c^2+c-1 (now we need to factorize) =(2c-1)(c+1) now put it back to cos(x) =(2cos(x)-1)(cos(x)+1) Hope that's better! [I have to be away from the keyboard for the next hour or so, sorry, but I'll be back later. If you still have problems and are in a rush, make a new post]